# Spin-orbit misalignment and residual eccentricity are evidence that neutron star-black hole mergers form through triple star evolution

### v1

@@ Abstract
There is growing evidence that a substantial fraction of the neutron star-black holes (NSBHs) detected through gravitational waves merge with non-zero eccentricity or large BH spin-orbit misalignment. This is in tension with the leading formation scenarios to date. Residual eccentricity rules out formation through isolated binary star evolution, while NS natal kicks and the unequal masses of NSBHs inhibit efficient pairing in dense stellar environments. Here, we report that all observed properties—NSBH merger rate, eccentricity, and spin-orbit misalignment—are explained by the high prevalence of massive stellar triples in the field. Modelling their evolution from the zero-age-main-sequence, we investigate NSBH mergers caused by gravitational perturbations from a tertiary companion. We show that the formation of the NS (its natal kick and the mass-loss in supernova, often reduced via binary envelope stripping) decisively impacts the triple stability, preferentially leaving behind surviving NSBHs in compact triple architectures. The rich three-body dynamics of compact, unequal-mass triples enables mergers across a wide range of orbital parameters without requiring fine-tuned highly inclined tertiary orbits and provides a natural explanation for an abundance of residual eccentricity and spin-orbit misalignment. We infer a total NSBH merger rate of $\mathcal{R}_{\rm NSBH}\sim 1$ – $23\,\rm Gpc^{-3}\,yr^{-1}$, with more than a few $10\,\%$ exhibiting residual eccentricity $e_{20}>0.1$ or large spin-orbit misalignment $\cos\theta_{\rm BH}<0$, consistent with current observations. Tertiary-driven NSBH mergers closely track the cosmic star formation rate due to their short delay times ($\sim 10$ – $100\,\rm Myr$), include a substantial fraction of burst-like highly eccentric systems ($e_{20}>0.9$), and almost universally retain eccentricities $e_{20}>10^{-3}$, potentially detectable by next-generation detectors. If evidence for eccentric and misaligned events solidifies, our results suggest that field triple dynamics is the dominant formation channel of gravitational-wave events of NSBH mergers.
@@ 1 Introduction
Three of the six potential neutron star-black hole (NSBH) mergers discovered through gravitational wave (GW) radiation (R. Abbott et al., 2020, 2021, 2023; A. G. Abac et al., 2024) exhibit enigmatic properties that challenge the standard formation scenario from the evolution of interacting isolated massive binary stars (e.g., M. Mapelli & N. Giacobbo, 2018; F. S. Broekgaarden et al., 2021; P. Drozda et al., 2022; Z. Xing et al., 2024). On the one hand, R. Abbott et al. (2021) report an 88 % probability that the black hole (BH) spin of GW200115 is misaligned to the binary orbital angular momentum vector by more than $90\,{\rm deg}$. Moreover, the BH spin of the NSBH candidate GW230529 shows strong evidence of being misaligned, but it cannot be ruled out that the event is a binary neutron star (NS) merger (A. G. Abac et al., 2024). If NSBH mergers result from the common-envelope evolution of isolated binary stars, G. Fragione et al. (2021) showed that such large spin-orbit tilts could only be accommodated with large NS natal kicks ($\gtrsim 150\,\rm km\,s^{-1}$) and high efficiencies for common-envelope ejection ($\alpha_{\rm CE}\gtrsim 3$). On the other hand, G. Morras et al. (2025) found evidence that the NSBH merger GW200105 (R. Abbott et al., 2021) retained a residual eccentricity $\gtrsim 0.1$ upon merging (excluding zero at more than 99 % confidence), which was later corroborated by M. de Lluc Planas et al. (2025). Any measurable residual eccentricity close to merger would imply that the binary must have had an extremely large eccentricity in the past because of highly efficient orbital circularisation by GW radiation (P. C. Peters, 1964). This would be wholly inconsistent with a formation from isolated massive binary stars (e.g., K. Belczynski et al., 2002; G. Fumagalli et al., 2024).
@@
Inferring a non-zero spin-orbit tilt and eccentricity in the aforementioned GW events is sensitive to assumptions of the parameter estimation. Regarding GW200115, I. Mandel & R. J. E. Smith (2021) use interacting binary star models to advocate for a BH spin prior limited to small magnitudes (as opposed to larger magnitudes allowed by R. Abbott et al. (2021)) leading to no support for significant spin-orbit misalignment. Meanwhile, G. Morras et al. (2025) find non-zero eccentricity of GW200105 for all of their chosen priors of the eccentricity except for log-uniform priors with a lower prior bound of $<10^{-4}$. Moreover, GW200105 was effectively a single-detector event recorded by LIGO Livingston while LIGO Hanford was not operational and Virgo yielded a low signal-to-noise ratio (R. Abbott et al., 2021). However, if evidence for large spin-orbit tilts and non-zero eccentricity becomes more robust in the growing population of detected NSBH mergers, it calls for alternative formation scenarios where the mergers are caused by some dynamical perturbation from the environment.
@@
In the literature, there exist four main formation channels in which NSBHs (as well as stellar binary BHs and binary NSs) could be dynamically driven to merge and which can generally recover a considerable fraction of non-zero eccentricity and spin-orbit misalignment. First, merging binaries can result from dynamical interactions in dense star clusters. However, multiple works found the rate of NSBH mergers in star clusters to be several orders of magnitude smaller than the estimated rate $\mathcal{R}_{\rm LVK}=55^{+127}_{-47}\,\rm Gpc^{-3}\,yr^{-1}$ (A. G. Abac et al., 2024) from GW detections (D. Clausen et al., 2013; Y.-B. Bae et al., 2014; C. Petrovich & F. Antonini, 2017; K. Belczynski et al., 2018; C. S. Ye et al., 2019; B.-M. Hoang et al., 2020; M. Arca Sedda, 2020; G. Fragione & S. Banerjee, 2020; C. S. Ye et al., 2020). This is because BHs and NSs are expected to spatially separate in clusters. The heavier BHs mass-segregate into the dense cluster core where their dynamical interactions act as a heating source (“BH burning”) that keeps the lighter NSs at larger cluster radii. Second, ultra-wide binaries (with semi-major axis $\gtrsim 10^{3}\,\rm AU$) can be driven to merger by torques exerted by their host galaxy and fly-bys of surrounding stars (E. Michaely & H. B. Perets, 2019; Y. Raveh et al., 2022; J. Stegmann et al., 2024). E. Michaely & S. Naoz (2022) found that the rate of NSBH mergers from wide binaries agrees with the observationally inferred rate if NSs and BHs receive zero natal kicks, which is uncertain. If NSs even receive moderate natal kicks $\gtrsim\mathcal{O}(1)\,\rm km\,s^{-1}$ the vast majority of the loosely bound ultra-wide binaries gets disrupted and the rate of NSBH mergers would be significantly suppressed (J. Stegmann et al., 2024). Third, NSBHs can merge in active galactic nuclei at high rates, but the details of this model are highly uncertain (Y. Yang et al., 2020; B. McKernan et al., 2020; H. Tagawa et al., 2021).
@@
In this work, we investigate the formation of NSBH mergers through the fourth channel, where compact object mergers are caused by the presence of a distant bound tertiary companion. This scenario is well-motivated by observations showing that most massive progenitor stars of NSs and BHs are found in hierarchical triples (M. Moe & R. Di Stefano, 2017). In these systems the gravitational perturbation from the tertiary companion can induce long-term, large-amplitude von Zeipel-Kozai-Lidov (ZKL) eccentricity oscillations of the inner binary (H. V. Zeipel, 1909; Y. Kozai, 1962; M. L. Lidov, 1962) and promote a GW-driven inspiral at close periapsis passage (E. B. Ford et al., 2000; L. Wen, 2003; F. Antonini et al., 2014; J. M. Antognini et al., 2014; F. Antonini et al., 2016, 2017; B. Liu et al., 2019; M. A. S. Martinez et al., 2020; G. Fragione & S. Banerjee, 2020; I. Bartos et al., 2023; A. Vigna-G\’omez et al., 2025). G. Fragione & A. Loeb (2019a, b) focussed on the formation of NSBH mergers from the evolution of the subset of triples with wide non-interacting inner binary stars (separated by $\geq 10\,\rm AU$). While they did find a significant fraction of tertiary-driven NSBH mergers with finite eccentricity and spin-orbit misalignment, the wide orbits of their triples get easily disrupted by NS natal kicks. Thus, the resulting merger rates are several orders of magnitude lower than the observationally inferred rate unless zero kicks are assumed. Similarly, A. A. Trani et al. (2022) considered the evolution of relatively wide compact object triples obtained from direct $N$-body simulations of dissolving young star clusters. They also recovered a considerable fraction of eccentric tertiary-driven NSBH mergers, albeit with a total merger rate that is one to three orders of magnitude lower than the observationally inferred range. In contrast, A. S. Hamers & T. A. Thompson (2019); J. Stegmann et al. (2022a) did include the high close binary fraction observed for massive stars (e.g., H. Sana et al., 2012) which leads to a much larger number of surviving systems and NSBH merger rates consistent with the detections. However, they did not fully model the three-body dynamics after NSBH formation nor investigated the detailed properties of any merger, leaving the characteristics of the bulk of NSBH mergers from triples largely unexplored and a comparison to GW data pending.
@@
Figure 1: Cartoon of hierarchical NSBH triples whose formation and evolution is studied in this work. The two inner binary stars form an NSBH with masses $m_{1}=m_{\rm BH}$ and $m_{2}=m_{\rm NS}$ which is orbited by a distant tertiary companion star with mass $m_{3}$. The inner and outer orbits are ellipses (denoted by indices “1” and “2”, respectively) that can be characterised by eccentricity vectors $\mathbf{e_{1}}$ and $\mathbf{e_{2}}$ that point towards the orbital periapses and have magnitudes equal to the orbital eccentricities $|\mathbf{e_{1}}|=e_{1}$ and $|\mathbf{e_{2}}|=e_{2}$, and by dimensionless orbital angular momentum vectors $\mathbf{j_{1}}$ and $\mathbf{j_{2}}$ which are perpendicular to the orbital planes (S. Tremaine et al., 2009). The two semi-major axes of the inner and outer orbits are hierarchical in the sense that the semi-major axes satisfy $a_{1}\ll a_{2}$. We also consider the evolution of the BH spin vector $\mathbf{S}_{\rm BH}$ and its angle $\theta_{\rm BH}$ with respect to $\mathbf{j_{1}}$.
@@
Figure 3: Parameter space of BHOB triples before the second SN. Blue contours show the density of triples with inner binaries that survive the second SN and form stable NSBH binaries. Red contours represent systems successfully forming stable NSBH triples. Scatter points highlight systems which will lead to tertiary-driven NSBH mergers with the colour indicating their residual eccentricity $e_{20}$ at a dominant GW frequency $f_{\rm GW}=20\,\rm Hz$ (L. Wen, 2003). The parameter window for surviving NSBH triples (red) can be understood as a result of several constraining limits, marked with hatched and dotted regions: Diagonally hatched and dotted regions indicate regions where the natal kick tends to disrupt the inner binary, the natal kick tends to disrupt the outer binary, and the SN mass-loss disrupts the outer binary, respectively (see Eqs. (7) – (9)). Below $a_{1}\approx 10\,\rm R_{\odot}$ inner binaries would undergo a stellar merger at ZAMS (horizontally hatched), and may only be occupied by systems surviving a common-envelope phase. The square hatched region indicates region of dynamical instability of the triple. The black dashed line shows $j_{\rm GW}=j_{\rm GR}$ for a fiducial triple with $m_{1}=10\,\rm M_{\odot}$, $m_{2}=1.5\,\rm M_{\odot}$, $m_{3}=1\,\rm M_{\odot}$, and $e_{2}=0.67$, respectively.
2 Methods
@@
We consider the evolution of hierarchical triples composed of an inner binary of two massive stars ($m_{1}\geq 10{\rm\,M_{\odot}}$, $m_{2}\geq 5{\rm\,M_{\odot}}$) and a less massive tertiary star ($m_{3}<5{\rm\,M_{\odot}}$). Our goal is to model the formation of stable triples that harbour an NSBH in the inner binary (“NSBH triple” hereafter, see Fig. 1) which is subsequently driven to merge by the gravitational perturbation from the tertiary companion. Because the stellar lifetime rapidly decreases with mass ($t\propto M^{-\alpha}$, where $\alpha\approx 2.5-3.5$), the NSBH will form before the tertiary leaves the Main Sequence (MS). Combined with negligible stellar winds of MS stars below $5{\rm\,M_{\odot}}$, this allows us to ignore the stellar evolution of the tertiary and treat it as a point mass. During the stellar evolution from the MS to the possible formation of NSBHs, we employ a model for binary star evolution on the inner orbit (described in Sec. 2.1), apply the effect of binary mass-loss and any supernova to the orbits of a triple (Sec. 2.2), and evolve the gravitational three-body dynamics after the NSBH has formed (Sec. 2.3). Not accounting for the gravitational three-body dynamics prior to the NSBH formation neglects the possibility that the inner binary star is driven to some interaction because of ZKL oscillations (A. S. Hamers & T. A. Thompson, 2019; S. Toonen et al., 2020; J. Stegmann et al., 2022a, b), although for most systems we expect them to be quenched by tides between the close inner binary stars.
2.1 Evolution of inner binaries

### v2

# Orbital Eccentricity And Spin-Orbit Misalignment Are Evidence That Neutron Star-Black Hole Mergers Form Through Triple Star Evolution

@@ Abstract
There is growing evidence that a substantial fraction of the neutron star-black holes (NSBHs) detected through gravitational waves merge with non-zero eccentricity or large BH spin-orbit misalignment. This is in tension with leading formation scenarios to date. Residual eccentricity rules out formation from isolated binary stars, while NS natal kicks and unequal masses of NSBHs inhibit efficient pairing in dense stellar environments. We report that all observed properties—NSBH merger rate, eccentricity, and spin-orbit misalignment—are explained by the high prevalence of massive stellar triples in the field. Modelling their evolution from the zero-age-main-sequence, we investigate NSBH mergers caused by gravitational perturbations from a tertiary companion. The NS formation decisively impacts the triple stability, preferentially leaving behind surviving NSBHs in compact triple architectures with mild hierarchies. The rich three-body dynamics of compact, unequal-mass triples enables mergers across a wide range of orbital parameters and provides a natural explanation for an abundance of residual eccentricity and spin-orbit misalignment. We infer a total NSBH merger rate of $\mathcal{R}_{\rm NSBH}\sim 1$$23\,\rm Gpc^{-3}\,yr^{-1}$ (within uncertainties on NS kicks) with more than a few $10\,\%$ exhibiting residual eccentricity $e_{20}>0.1$ or large spin-orbit misalignment $\cos\theta_{\rm BH}<0$, consistent with current observations. The mergers closely track the cosmic star formation rate due to short delay times ($\sim 10$$100\,\rm Myr$), include a substantial fraction of burst-like highly eccentric systems ($e_{20}>0.9$), and almost universally retain eccentricities $e_{20}>10^{-3}$ detectable by next-generation detectors. If evidence for eccentric and misaligned events solidifies, our results suggest that triple dynamics is the dominant formation channel of gravitational-wave events from NSBH mergers.
@@ I Introduction
Three of the seven potential neutron star-black hole (NSBH) mergers discovered through gravitational wave (GW) radiation (Abbott et al., 2020, 2021, 2023; Abac et al., 2024; The LIGO Scientific Collaboration et al., 2025b) exhibit enigmatic properties that challenge the standard formation scenario from the evolution of interacting isolated massive binary stars (e.g., Mapelli and Giacobbo, 2018; Broekgaarden et al., 2021; Drozda et al., 2022; Xing et al., 2024). On the one hand, Abbott et al. (2021) report an 88 % probability that the black hole (BH) spin of GW200115 is misaligned to the binary orbital angular momentum vector by more than $90\,{\rm deg}$. Moreover, the BH spin of the NSBH candidate GW230529 shows strong evidence of being misaligned, but it cannot be ruled out that the event is a binary neutron star (NS) merger (Abac et al., 2024). If NSBH mergers result from the common-envelope evolution of isolated binary stars, Fragione et al. (2021) showed that such large spin-orbit tilts could only be accommodated with large NS natal kicks ($\gtrsim 150\,\rm km\,s^{-1}$) and high efficiencies for common-envelope ejection ($\alpha_{\rm CE}\gtrsim 3$). On the other hand, Morras et al. (2025) found evidence that the NSBH merger GW200105 (Abbott et al., 2021) retained a residual eccentricity $\gtrsim 0.1$ upon merging (excluding zero at more than 99 % confidence), which aligns with earlier findings in the context of tests for alternative theories of gravity (Fei and Yang, 2024) and which was later corroborated by de Lluc Planas et al. (2025), Kacanja et al. (2025), and Jan et al. (2025). Any measurable residual eccentricity close to merger would imply that the binary must have had an extremely large eccentricity in the past because of highly efficient orbital circularisation by GW radiation (Peters, 1964). This would be wholly inconsistent with a formation from isolated massive binary stars (e.g., Belczynski et al., 2002; Fumagalli et al., 2024).
@@
Inferring a non-zero spin-orbit tilt and eccentricity in the aforementioned GW events is sensitive to assumptions of the parameter estimation. Regarding GW200115, Mandel and Smith (2021) use interacting binary star models to advocate for a BH spin prior limited to small magnitudes (as opposed to larger magnitudes allowed by Abbott et al. (2021)) leading to no support for significant spin-orbit misalignment. Meanwhile, Morras et al. (2025) find non-zero eccentricity of GW200105 for all of their chosen priors of the eccentricity except for log-uniform priors with a lower prior bound of $10^{-4}$. Moreover, GW200105 was effectively a single-detector event recorded by LIGO Livingston while LIGO Hanford was not operational and Virgo yielded a low signal-to-noise ratio (Abbott et al., 2021). However, if evidence for large spin-orbit tilts and non-zero eccentricity becomes more robust in the growing population of detected NSBH mergers, it calls for alternative formation scenarios where the mergers are caused by some dynamical perturbation from the environment.
@@
In the literature, there exist four main formation channels in which NSBHs (as well as stellar binary BHs and binary NSs) could be dynamically driven to merge and which can generally recover a considerable fraction of non-zero eccentricity and spin-orbit misalignment. First, merging binaries can result from dynamical interactions in dense star clusters. However, multiple works found the rate of NSBH mergers in star clusters to be several orders of magnitude smaller than the estimated rate $\mathcal{R}_{\rm LVK}=9.1$ – $84\,\rm Gpc^{-3}\,yr^{-1}$ (The LIGO Scientific Collaboration et al., 2025a) from GW detections (Clausen et al., 2013; Bae et al., 2014; Petrovich and Antonini, 2017; Belczynski et al., 2018; Ye et al., 2019; Hoang et al., 2020; Arca Sedda, 2020; Fragione and Banerjee, 2020; Ye et al., 2020). This is because BHs and NSs are expected to spatially separate in clusters. The heavier BHs mass-segregate into the dense cluster core where their dynamical interactions act as a heating source (“BH burning”) that keeps the lighter NSs at larger cluster radii. Second, ultra-wide binaries (with semi-major axis $\gtrsim 10^{3}\,\rm AU$) can be driven to merger by torques exerted by their host galaxy and fly-bys of surrounding stars (Michaely and Perets, 2019; Raveh et al., 2022; Stegmann et al., 2024). Michaely and Naoz (2022) found that the rate of NSBH mergers from wide binaries agrees with the observationally inferred rate if NSs and BHs receive zero natal kicks, which is uncertain. If NSs even receive moderate natal kicks $\gtrsim\mathcal{O}(1)\,\rm km\,s^{-1}$ the vast majority of the loosely bound ultra-wide binaries gets disrupted and the rate of NSBH mergers would be significantly suppressed (Stegmann et al., 2024). Third, NSBHs can merge in active galactic nuclei at high rates, but the details of this model are highly uncertain (Yang et al., 2020; McKernan et al., 2020; Tagawa et al., 2021).
@@
In this work, we investigate the formation of NSBH mergers through the fourth channel, where compact object mergers are caused by the presence of a distant bound tertiary companion. This scenario is well-motivated by observations showing that most massive progenitor stars of NSs and BHs are found in hierarchical triples (Moe and Di Stefano, 2017). In these systems the gravitational perturbation from the tertiary companion can induce long-term, large-amplitude von Zeipel-Kozai-Lidov (ZKL) eccentricity oscillations of the inner binary (Zeipel, 1909; Kozai, 1962; Lidov, 1962) and promote a GW-driven inspiral at close periapsis passage (Ford et al., 2000; Wen, 2003; Antonini et al., 2014; Antognini et al., 2014; Antonini et al., 2016, 2017; Liu et al., 2019; Martinez et al., 2020; Fragione and Banerjee, 2020; Bartos et al., 2023; Vigna-G\’omez et al., 2025). Fragione and Loeb (2019b, a) focussed on the formation of NSBH mergers from the evolution of the subset of triples with wide non-interacting inner binary stars (separated by $\geq 10\,\rm AU$). While they did find a significant fraction of tertiary-driven NSBH mergers with finite eccentricity and spin-orbit misalignment, the wide orbits of their triples get easily disrupted by NS natal kicks. Thus, the resulting merger rates are several orders of magnitude lower than the observationally inferred rate unless zero kicks are assumed. Similarly, Trani et al. (2022) considered the evolution of relatively wide compact object triples obtained from direct $N$-body simulations of dissolving young star clusters. They also recovered a considerable fraction of eccentric tertiary-driven NSBH mergers, albeit with a total merger rate that is one to three orders of magnitude lower than the observationally inferred range. In contrast, Hamers and Thompson (2019); Stegmann et al. (2022a) did include the high close binary fraction observed for massive stars (e.g., Sana et al., 2012) which leads to a much larger number of surviving systems and NSBH merger rates consistent with the detections. However, they did not fully model the three-body dynamics after NSBH formation nor investigated the detailed properties of any merger, leaving the characteristics of the bulk of NSBH mergers from triples largely unexplored and a comparison to GW data pending.
@@
Figure 1: Cartoon of hierarchical NSBH triples whose formation and evolution is studied in this work. The two inner binary stars form an NSBH with masses $m_{1}=m_{\rm BH}$ and $m_{2}=m_{\rm NS}$ which is orbited by a distant tertiary companion star with mass $m_{3}$. The inner and outer orbits are ellipses (denoted by indices “1” and “2”, respectively) that can be characterised by eccentricity vectors $\mathbf{e_{1}}$ and $\mathbf{e_{2}}$ that point towards the orbital periapses and have magnitudes equal to the orbital eccentricities $|\mathbf{e_{1}}|=e_{1}$ and $|\mathbf{e_{2}}|=e_{2}$, and by dimensionless orbital angular momentum vectors $\mathbf{j_{1}}$ and $\mathbf{j_{2}}$ which are perpendicular to the orbital planes (Tremaine et al., 2009). The two semi-major axes of the inner and outer orbits are hierarchical in the sense that the semi-major axes satisfy $a_{1}\ll a_{2}$. We also consider the evolution of the BH spin vector $\mathbf{S}_{\rm BH}$ and its angle $\theta_{\rm BH}$ with respect to $\mathbf{j_{1}}$.
@@
Figure 3: Parameter space of BHOB triples before the second SN. Blue contours show the density of triples with inner binaries that survive the second SN and form stable NSBH binaries. Red contours represent systems successfully forming stable NSBH triples. Scatter points highlight systems which will lead to tertiary-driven NSBH mergers with the colour indicating their residual eccentricity $e_{20}$ at a dominant GW frequency $f_{\rm GW}=20\,\rm Hz$ (Wen, 2003). The parameter window for surviving NSBH triples (red) can be understood as a result of several constraining limits, marked with hatched and dotted regions: Diagonally hatched and dotted regions indicate regions where the natal kick tends to disrupt the inner binary, the natal kick tends to disrupt the outer binary, and the SN mass-loss disrupts the outer binary, respectively (see Eqs. (7) – (9)). Below $a_{1}\approx 10\,\rm R_{\odot}$ inner binaries would undergo a stellar merger at ZAMS (horizontally hatched), and may only be occupied by systems surviving a common-envelope phase. The square hatched region indicates region of dynamical instability of the triple. The black dashed line shows $j_{\rm GW}=j_{\rm GR}$ for a fiducial triple with $m_{1}=10\,\rm M_{\odot}$, $m_{2}=1.5\,\rm M_{\odot}$, $m_{3}=1\,\rm M_{\odot}$, and $e_{2}=0.67$, respectively.
II Methods
@@
We consider the evolution of hierarchical triples composed of an inner binary of two massive stars ($m_{1}\geq 10{\rm\,M_{\odot}}$, $m_{2}\geq 5{\rm\,M_{\odot}}$) and a less massive tertiary star ($m_{3}<5{\rm\,M_{\odot}}$). Our goal is to model the formation of stable triples that harbour an NSBH in the inner binary (“NSBH triple” hereafter, see Fig. 1) which is subsequently driven to merge by the gravitational perturbation from the tertiary companion. Because the stellar lifetime rapidly decreases with mass ($t\propto M^{-\alpha}$, where $\alpha\approx 2.5-3.5$), the NSBH will form before the tertiary leaves the Main Sequence (MS). Combined with negligible stellar winds of MS stars below $5{\rm\,M_{\odot}}$, this allows us to ignore the stellar evolution of the tertiary and treat it as a point mass. During the stellar evolution from the MS to the possible formation of NSBHs, we employ a model for binary star evolution on the inner orbit (described in Sec. II.1), apply the effect of binary mass-loss and any supernova to the orbits of a triple (Sec. II.2), and evolve the gravitational three-body dynamics after the NSBH has formed (Sec. II.3). Not accounting for the gravitational three-body dynamics prior to the NSBH formation neglects the possibility that the inner binary star is driven to some interaction because of ZKL oscillations (Hamers and Thompson, 2019; Toonen et al., 2020; Stegmann et al., 2022a, b), although for most systems we expect them to be quenched by tides between the close inner binary stars.
II.1 Evolution of inner binaries